OpenStudy (anonymous):

Write a rule for the function represented by the table. x y 0 13 1 28 2 43 3 58 (1 point) • y = 15 + 13x • y = 14 + 12x • y = 13 + 15x • y = 12 + 14x my answer is A check my work plz

OpenStudy (tkhunny):

Try x = 0 and see which gives y = 13.

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

which school

OpenStudy (tkhunny):

Help with what? Did you try x = 0, yet?

OpenStudy (anonymous):

idk how to do it

OpenStudy (tkhunny):

I do not believe that. y = 15 + 13x What happens when x = 0? y = 15 + 13(0) = 15 + 0 = 15 For the first one, when x = 0, then y = 15. Is that consistent with the chart? Do the rest.